Let be H - the point of intersection of the heights of an acute-angled triangle ABC. From points A and C drawn tangents AK and CT to the circle drawn on the line segment BH as in diameter. Let be 15 and 17 - the lengths of these tangents. What is the smallest possible side length AC? In response, write down the square of the length AC.
1
Expert's answer
2020-12-17T15:03:42-0500

AK2 + AC2 = CT2

152 + AC2 = 172

AC2 = 289 - 225

AC2 = 64

AC = 8

AC2 = 64